OpenStudy (anonymous):

Use spherical coordinates to evaluate the triple integral(x^2 + y^2 + z^2) dV, where E is the ball:x^2 + y^2 + z^2 < or =81.

OpenStudy (anonymous):

Doesn't x^2 + y^2 + z^2 = p^2? Then I'd just have p^4 sin.phi drho dphi dtheta

OpenStudy (anonymous):

yes. So you integral will be: \(\int_0^{81}\int_{-\pi/2}^{\pi/2 }\int_0^{2\pi} p^2 d\theta d\phi dp\)

OpenStudy (anonymous):

∫x2 u2 z2 d x = 1/3 x3 u2 z2

OpenStudy (experimentx):

\[ \int_0^9 \int_{0}^{\pi } \int_0^{2\pi} r^4 \sin \theta d\phi d\theta dr \]

OpenStudy (anonymous):

Would my limits of integration be: 0-p-9, 0-theta-2pi, 0-phi-pi? and then would I still have to put my \[\rho^2 \sin \phi\] before the differentials?

OpenStudy (experimentx):

yes ... looks like you you took theta for phi. usually phi is used for azimuthal coordinate.

OpenStudy (anonymous):

I usually use phi as the angle of declination. Do you have a math background or physics background? Just curious

OpenStudy (experimentx):

physics

OpenStudy (anonymous):

Ahh, I heard that physics people flip theta and phi around for some pointless reason. :)

OpenStudy (experimentx):

for the sake of ease ...